OpenStudy (anonymous):

graph y=-3/4x^3 with five points one point with x=0 , two points with negative x values and two points with positive x values

OpenStudy (raffle_snaffle):

make a table and follow the directions

OpenStudy (raffle_snaffle):



OpenStudy (anonymous):

okay, but what points do I use?

OpenStudy (raffle_snaffle):

Any points you want

OpenStudy (raffle_snaffle):

two numbers below zero and two above zero.

OpenStudy (anonymous):

like -1, -2, 1,2

OpenStudy (raffle_snaffle):

that is fine

OpenStudy (anonymous):

then what do I put for y?

OpenStudy (anonymous):

http://prntscr.com/3c2s72

OpenStudy (anonymous):

you solve the function

OpenStudy (anonymous):

to find the y for certain x

OpenStudy (anonymous):

I don't get it. I made a table and put for my x values 0,-1,-2,1,2 now what do I put for my y values??

OpenStudy (anonymous):

y=-3/4x^3 (1) at x you put the random numbers like x=-2,-1,0,1,2 and you solve (1) to find the respective y's for example: x=0: y=-3/4(0)^3=0 u do the same with the other

OpenStudy (anonymous):

okay hold on ill try to solve them

OpenStudy (raffle_snaffle):

@ jovel I think the guy know how to plug x values in a function and obtain an output value.

OpenStudy (anonymous):

I got ( 0,0) (-1,0.75),(-2,6) (1,-0.75)(2,-6) correct?

OpenStudy (anonymous):

uhhh @raffle_snaffle read above I'm not so sure...

OpenStudy (anonymous):

it's correct @jbrocksmysocks96 (y)

OpenStudy (anonymous):

thank you so much

OpenStudy (raffle_snaffle):

thanks @jovel